OpenStudy (anonymous):

What is the simplified form of (x+7)/(x+3) + (x-4)/(3)?

OpenStudy (anonymous):

So it basically looks like this \[\frac{ (x+7) }{ (x+3) } + \frac{ (x-4) }{ 3 }\]

OpenStudy (anonymous):

I'm just confused what I need to multiply the denoms by? Like do I multiply the first equation by 3 and the second one by (x+3)?

OpenStudy (camzzzie):

just go on wolframalpha and search the problem

OpenStudy (anonymous):

uhh sorry, I wanna know how to actually solve it.

OpenStudy (bibby):

we want to get them into a common denominator like so: \(\large \dfrac{ (x+7) }{ (x+3) } + \dfrac{ (x-4) }{ 3 }=\dfrac{?}{(3)(x+3)}\) we can do this by multiplying the left term by 3 and the right term by (x+3) \(\large \dfrac{ (x+7) }{ (x+3) } + \dfrac{ (x-4) }{ 3 }=\dfrac{(3)(x+7)+(x+3)(x-4)}{(3)(x+3)}\)

OpenStudy (anonymous):

ok makes sense. is it ok if I evaluate and you tell me if I'm right?

OpenStudy (bibby):

we then expand: